Write the answer in scientific notation. 1.4x10^-4/5.6x10^-7
Answer by drk(1908)

step 1 separate the two as
(i) (1.4 x 10^(-4)) / (5.6 x 10^(-7))
step 2 - divide (reduce the front numbers to get
1.4 / 5.6 = 1/4 = .25
step 3 - use rules of exponents to divide the bases. We get
10^(-4) / 10^(-7) = 10^(3)
step 4 - put steps 2 and 3 together to get
(ii) .25 x 10^3
step 5 - .25 is not in scientific notation, so we adjust it to get
.25 = 2.5 x 10^(-1)
step 6 - put this back into (ii) to get
(iii) 2.5 x 10^2